OpenStudy (anonymous):

help me with this log[5]√125

Parth (parthkohli):

\[\large \log_5 \sqrt{125} = \log_5 5^{3/2} = \frac{3}{2} \log_5 5 = \cdots\]

OpenStudy (anonymous):

@ParthKohli sorry i don't get it

OpenStudy (yttrium):

@bless just remember that \[\log_{a}a = 1 \] in your case \[\frac{ 3 }{ 2 }\log_{5}5 = x \] Hence, x = ?

Parth (parthkohli):

Sorry, I am back. Can we write \(\sqrt{125} = \sqrt{5^3} = 5^{3/2}\)?

OpenStudy (anonymous):

@ParthKohli yes

Parth (parthkohli):

That is what I have done here. Also note that \(\log\left(a^m\right) = m \log(a)\)

OpenStudy (anonymous):

why did √5^3 turns into 5^3/2?

OpenStudy (anonymous):

@ParthKohli ?